Find and correct the errors in the following mathematical statements.
(3x + 2)^2 = 3x^2 +6x+4 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = (3x+2)^2= (3x)^2+2^2+2x2x3x = 9x^2+4+12x≠ RHS
The correct statement is (3x + 2)^2= 9x^2+4+12x

Thanks